OpenStudy (anonymous):

suppose a curve is given by the parametric equations: x=2+cost, y=3+sint as t varies in the interval [0,2pi] make a table of values for the curve using t=0, /6, /4, /3, /2 ,3/2, 2

OpenStudy (anonymous):

Not too difficult. What you have to do is construct a table like so:  Now in the columns for x and y, you must evaluate x = 2 + cost, and y = 3 + sint for the given values of t in the left-most column. For example, x(0) = 2 + cos(0) = 2 + 1 = 3 y(0) = 3 + sin(0) = 3 + 0 = 3. And so on.